Draw the that z 0 m at t = 08.(Figuro 1) graph for the velocity 1 o1 No elements selected r(m) y, (m/s) t (s) 10 12 Can somone please explain how to graph this, please show work and explain where you get the numbers. Thank you
Answer #1

From the given velocity vs. time graph, we can see that, for the first 4.00 s, velocity is constant.

so, the slope of the position time curve should be equal to -3.00.

v = -12 / 4 = -3 m/s

Now velocity is increased to +3 m/s/.

v = -6 - (-12) / (6 - 4) = +3 m/s.

From 6 to 12 s, again velocity is constant . so, position vs. time graph is a straight line.

We can see that displacement of the object is 0.